• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 00:25:41 on October13,2005

    Entry number 156686

    keyword=cranked HV on Left Horiz.

    Increased HV to 2100 (increase of 200V) on Left Horizontal segment.
    See cosmics in scalers now (1 to 3 Hz).

    For 0.2 uA we see rates on L-arm
    Vert. Segment 12 kHz
    Horiz. Segment 1.3 kHz (1/10th)
    On R-arm the 2 segments are more balanced, about 75 & 83 kHz

    Crank HV on L-arm Horiz another 100V to 2200V.
    Vert 13 Hz
    Horiz 2.2 kHz

    Run 2817 with dP = -2% on L-arm, but dP = 0 on R-arm.
    P_L = 3.0654 still.
    P_R = 3.1776 (new optimum)

    Wrong prescales on T2. Run 2818 is a repeat with better PS.

    Now its clear the L-arm Horiz segment makes a signal (fig 1)
    but the amplitude is way out of whack (1/5 th of other 3 PMTs).
    David L. wants to take access to remove a filter.

    We could probably do the alignment with this unbalanced rate
    (would need a slightly more careful offline analysis to avoid
    bias from the dominant detector) but we decide its best to
    verify we understand what the detector is doing ...

    I (Bob) am leaving now. Rob F. to take over this part.


    FIGURE 1